Recent results from the BES collaboration

Zhengguo Zhao, for the BES Collaboration Institute of High Energy Physics of Chinese Acadmy of Sciences Beijing 100039, China

The status of the upgraded BEijing Spectrometer(BESII) and the measurement of the R values in the center-of-mass (cm) energy region of 2-5 GeV done with BESII are reported. Recent results based on the analysis of the data collected with BESI are summarized.

1. INTRODUCTION

The BEijing Spectrometer (BES) at Beijing Electron Positron Collider (BEPC), so far the only e^{+}e^{−} machine in the world which covers the cm energy in 2-5 GeV and directly produce charmonium and charmed mesons, has been running for 10 years. Both BEPC and BES were upgraded from 1995 to 1997. The upgrade of BEPC included moving the insertion quadrupoles closer to the interaction point and increasing the total RF voltage. The bunch length was shortened, and the vertical beta function at the interaction point was reduced. There is a factor of 1.5-2 improvement in the luminosity. The hadronic event rate recorded at J/ψ peak can now reach 6-8 Hz.

The BES is a conventional cylindrical magnetic detector and is described in detail in Ref. [1]. The upgrades of BES [2] included replacement of the central drift chamber with a vertex chamber composed of 12 tracking layers. The vertex chamber was rebuilt from the Mark III endplates and beryllium beam pipe and provides a spatial resolution of about 90 μm. A new barrel time-of-flight counter with shortened light pipes improves on the BES I design and provides 180 ps resolution. A new main drift chamber replaces the aging original, providing a similar momentum and dE/dx resolution to the old one. With the upgraded data acquisition system, the dead time has been reduced from 20 ms to 10 ms per event. Table 1 compares the performance of BESI with that of BESII.

2. R SCAN WITH BESII

QED running coupling constant α(s) and anomalous magnetic moment a_{μ} = (g − 2)/2 of the μ are fundamental physics observables to test the Standard Model [3]. Their theoretical precisions so far are limited by the second order loop effects from the hadronic vacuum polarization. About 50% and 20% of the errors in α(M_{Z}^{2}) and a_{μ}, respectively, are due to the uncertainty of the values of R = σ(e^{+}e^{−} → hadrons)/σ(e^{+}e^{−} → μ^{+}μ^{−}) in the energy below 5 GeV, where the denominator is the lowest order QED cross section, σ(e^{+}e^{−} → μ^{+}μ^{−}) = σ_{μμ}^{0} = 4πα^{2}/3s. To reduce the uncertainties in the values of R in 2-5 GeV region, we scanned 6 points covering cm energy from 2.6 to 5.0 GeV in 1998 (Run I) [4], and 85 points distributed in 2.0-4.8 GeV energy region with BESII in Table 1 Comparison of the major parameters with BESI and BESII

System  Parameter  BESI  BESII
VC  σ_{xy}(μ)  250 (CDC)  90
planes  4  12
MDC  σ_{xy}(μ)  200-250  190-210
Δp/p(%)(p in GeV) 1.76\sqrt{(1 + p^{2})} 1.78\sqrt{(1 + p^{2})}
TOF  σ_{t}(ps)  370  180
L_{attenuation}(m)  1-1.2  3.5-5.5
BSC ΔE/E(%)(E in GeV)  25  21
ESC ΔE/E(%)(E in GeV)  23  22
DAQ dead time/event(ms)  20  10

1999(Run II) [5]. The R values obtained at the six energy points are shown in Table 2. The preliminary R values at 9 points and the online R values for most of the points from Run II are illustrated in Fig. 1. The R scan done with BESII at BEPC can reduce the present uncertainties in R from ∼15% to ∼7% for E_{cm} < 4.0 GeV and ∼10% for 4.0 GeV < E_{cm} < 5.0 GeV. We expect to have the preliminary result for all the points scanned in the Summer of 2000.

Table 2 R values measured with BESII from Run I.

E_{cm}(GeV) 2.6 3.2 3.4 3.55 4.6 5.00
R  2.64 2.21 2.38 2.23 3.58 3.47
Stat. error 0.05 0.07 0.07 0.06 0.20 0.32
Sys. error 0.19 0.13 0.16 0.16 0.29 0.29

3. RESULTS ON CHARMONIUM

Based on the data collected with BESI shown in Table 3, BES has been studying the light hadron spectroscopy and searching for glueball, studying the properties of the charmonium and charmed mesons from the J/ψ, ψ(2S) and the charmed meson decay.

3.1. Results from ψ(2S) data

• J/ψ → l^{+}l^{−}braching fraction [6] The branching fraction for the leptonic decays J/ψ → e^{+}e^{−} (B_{e}) and μ^{+}μ^{−} (B_{μ}), the

basic parameters of the J/ψ resonance, can be used to determine the total number of J/ψ events in a wide variety of measurements that take advantage of the clean experimental J/ψ → l^{+}l^{−}(l = e or μ)

Figure 1. The online R values (triangle) in the energy region below 5 GeV. Notice that the background, the radiative correction and the detection efficiency have not been taken into account. The solid sqaures represent the preliminary R-values obtained from BESII data, of which the total errors are all conservatively assigned to be 10%. The final R-values of the 6 points listed in table 2 are also plotted with solid dots for comparison.

We determined J/ψ leptonic branching fraction from a comparison of the exclusive and inclusive processes: ψ′ → π^{+}π^{−}J/ψ with J/ψ → l^{+}l^{−}+anything.

The branching fractions are measured to be B_{e} = (5.90 ± 0.05 ± 0.10)% and B_{μ} = (5.84 ± 0.06 ± 0.10)%, which leads to B_{e}/B_{μ} = 1.011 ± 0.013 ± 0.016.

Assuming B_{e} = B_{μ}, we find a combined leptonic branching fraction of B_{l} = B(J/ψ → l^{+}l^{−}) = (5.87 ± 0.04 ± 0.09)% and obtain a new world average B_{l} = (5.894 ± 0.086)%, of which the error is about half that in the 1998 PDG.

Table 3 Summary of the data taken with BESI.

E_{cm}(GeV) Sample size
J/ψ resonance 7.8 × 10^{6}
ψ(2S)  3.7 × 10^{6}
4.03  22.3pb^{−}1

• Hadronic ψ(2S) decay BES has measured a large number of ψ(2S) decay branching fractions. Many for the

first time. Table 4 lists the recent preliminary results [7].

• χ_{cJ} hadronic decay Besides about 30 channels of χ_{cJ} hadronic decay, of which many are for the first time and

many improved PDG values, BES measured χ_{cJ} branching ratios of some new hadronic decay channels, as shown in Table 5.

Table 4 Hadronic ψ(2S) decay branching fraction measured with BES

Channels Br_{P DG}(×10^{−4}) Br_{BES}(×10^{−4}) Q_{h}^{BES} = \frac{Br_{ψ′}}{Br_{J/ψ}} (%)
ωη  –  < 0.33  < 2.1
π^{+}π^{−}π^{0}  0.9±0.5  1.06 ± 0.11 ± 0.16  7.1 ± 1.6
K\overline{K}π  –  1.25 ± 0.18 ± 0.26  2.1 ± 0.6
γf_{2}(1270)  –  2.36 ± 0.24 ± 0.32  17.1 ± 2.9
γf_{J}(1710)  –  1.24 ± 0.20 ± 0.16  12.8 ± 3.4
ϕK^{+}K^{−}  –  0.61 ± 0.18 ± 0.15  7.3 ± 3.0
ϕπ^{+}π^{−}  –  1.76 ± 0.22 ±0.24  22.0 ± 5.3
ϕf_{0}(980)  –  0.66 ±0.17 ± 0.07  20.5 ± 8.1
ωπ^{+}π^{−}  –  4.47 ±0.57 ± 0.54  6.3 ± 1.4
ωK^{+}K^{−}  –  1.26 ±0.43 ± 0.39  17.0 ± 9.6
ωpp̄  –  0.60 ±0.23 ± 0.14  4.6 ± 2.2
ϕpp̄  –  0.86 ±0.50 ± 0.20  18.9 ± 13.4
π^{+}π^{−}π^{0}pp̄  –  3.85 ±0.37 ± 0.56  16.7±7.1
ηπ^{+}π^{−}pp̄  –  2.58 ±0.68 ± 0.73
ηpp̄  –  0.84 ±0.48 ± 0.48  4.00±4.00
pp̄  1.9 ± 0.5  2.62 ± 0.17 ± 0.55  12.3 ± 2.8
Λ\overline{Λ}  < 4.0  1.89 ±0.21 ± 0.21  14.0 ± 2.7
Σ^{0}\overline{Σ}^{0}  –  1.20 ±0.40 ± 0.30  9.2 ± 4.1
Ξ^{−}\overline{Ξ}^{+}  < 2.0  1.00 ±0.30 ± 0.10  11.1 ± 4.3
Δ^{++}\overline{Δ}^{−−}  –  1.34 ±0.11 ± 0.33  12.2 ± 4.5
Σ^{∗+}\overline{Σ}^{∗−}  –  1.10 ±0.30 ± 0.30  10.7 ± 4.3

Table 5 Preliminary branching ratios of χ_{cJ} hadronic decay.

decay channels  BES (×10^{−3})  PDG(×10^{−3})
Br(χ_{c0} → π^{0}π^{0}) 2.80 ± 0.32 ± 0.51  3.1 ± 0.6
Br(χ_{c2} → π^{0}π^{0}) 0.92 ± 0.27 ± 0.52  1.1 ± 0.3
Br(χ_{c0} → ηη) 2.03 ± 0.84 ± 0.58  2.5 ± 1.1
Br(χ_{c2} → ηη)  < 2.5  0.8 ± 0.5

3.2. Study of J/ψ decay

• Partial Wave Analysis of J/ψ → γK^{+}K^{−}π^{+}π^{−} [8] For J/ψ → γK^{+}K^{−}π^{+}π^{−}, we found that the K^{∗}K̄^{∗} contribution peaks strongly near

threshold, which can be fitted with a broad 0^{−+} resonance with mass M = 1800 ± 100 MeV, width Γ = 500 ± 200 MeV. A broad 2^{++} resonance peaking at 2020 MeV is also required with width ∼ 500 MeV. There is further evidence for a 2^{−+} component peaking at 2.55 GeV. The non-K^{∗}K̄^{∗} contribution is close to phase space; it peaks at 2.6 GeV and is very different from K^{∗}K̄^{∗}.

• Measurement of m_{η_{c}} and Γ_{η_{c}} The decay modes of η_{c}: K^{+}K^{−}π^{+}π^{−}, π^{+}π^{−}π^{+}π^{−}, K^{±}K_{S}^{0}π^{∓}(with K_{S}^{0} → π^{+}π^{−}), ϕϕ(with

ϕϕ → K^{+}K^{−}K^{+}K^{−}) and K^{+}K^{−}π^{0} from J/ψ → γη_{c} have been analyzed and the mass of η_{c} is measured to be 2976.6 ± 2.9 ± 1.3 MeV.

Combining the results from BES data analysis of ψ(2S) → γη_{c}, m_{η_{c}} = 2976.3±2.3±1.2 MeV and Γ_{η_{c}} = 11.0 ± 8.1 ± 4.1 MeV are obtained [9].

• The analysis of J/ψ → γγV (V = ρ, ϕ)

We studied the J/ψ → γγV (V = ρ, ϕ) decay channels and found the evidence of the η(1430) in both decay channels. The mass and width of the η(1430) resonance are found to be 1431±17 MeV and 88±28 MeV from γγρ mode and 1427±13 MeV and 77±46 MeV from γγϕ mode [10]. In order to understand the structure of η(1430), other J/ψ decay channels should be analyzed.

• N^{∗} results from J/ψ decay [11] The behave of N^{∗} baryon in J/ψ decay to baryon-antibaryon final states provides us a

laboratory for the study of N^{∗} baryon, especially in the mass range of ∼1-2 GeV. Using PWA, we have studied the J/ψ → pp̄η and pp̄π^{0} decay. Two resonances with J^{P} = \frac{1}{2}^{−} are observed. Their masses and widths are: M = 1540_{−17}^{+15} MeV with Γ = 178_{−22}^{+20} MeV; and M = 1648_{−16}^{+18} MeV with Γ = 150 MeV respectively. These two states are considered to be the nucleon resonances S_{11}(1535) and S_{11}(1650).

3.3. Charmed mesons

Though many precise measurements of exclusive decay branching fractions of charmed mesons has been made to understand the heavy flavor physics, our knowledge on the inclusive decay properties of the charmed mesons is still limited by the experimental data. Using 22.3 pb^{−1} data collected at 4.03 GeV, we directly measured the inclusive \overline{K}^{∗0} and K^{∗0} decay branching fractions of the D_{s}^{+} meson. From 183.6 ± 24 ± 6.5 D_{s}^{+}D_{s}^{−} pairs and 7.5±4.4 D_{s}^{+} → \overline{K}^{∗0}X observed events, we deduced the results of: Br(D_{s}^{+} → \overline{K}^{∗0}X) = (20.3 ± 11.9 ± 1.8)% and Br(D_{s}^{+} → K^{∗0}X) < 7.5% at 90% C.L. [12]. Our results may indicate that some decay of D_{s}^{+} to \overline{K}^{∗0} are still missing.

4. PLANS FOR THE NEAR FUTURE OF BESII

The plans for our near future are:

• continue analyzing the R scan data and have the final R values for all the energy points in the summer of 2000;

• continue analyzing the data taken with BESI;

• collect ∼ 50 × 10^{6} J/ψ events with BESII within 3 years.

With this amount of J/ψ event sample, ∼6 times as more data as the largest sample existing in the world, BES collaboration can systematically study the light hadron spectroscopy and N^{∗} baryons; search for glueball and hybrid states; as well as explore the rare decays in the tau charm sector.

The results presented here are preliminary and the results based on J/ψ data are from only the China side of the BES collaboration.

We would like to thank BEPC staffs and IHEP computing center for their great contribution made to the R scan. We also wish to acknowledge useful discussions with B. Andersson, H. Burkhardt, M. Davier, B. Pietrzyk, T. Sjöstrand, M. L. Swartz, J. M. Wu, C. X. Zhang, X. M. Zhang, and G. D. Zhao. We especially thank M. Tigner for major contributions not only to BES but also to the operation of the BEPC. This work is supported in part by the National Natural Science Foundation of China under Contract Nos. 19290400, 19991480, and 19825116; the Chinese Academy of Sciences under contract No. KJ95T-03(IHEP); and by the Department of Energy under Contract Nos. DE-FG03-93ER40788 (Colorado State University), DE-AC03-76SF00515 (SLAC), DE-FG03-94ER40833 (U Hawaii), DE-FG03-95ER40925 (UT Dallas).

REFERENCES

1. J.Z. Bai et al., (BES Collab.), Nucl. Instrum. Methods 344(1994)319.

2. J. Li, ”The Upgraded BES Detector (BESII)” talk given at Workshop on Relativistic Heavy Ion Collisions and Quark Matter Physics, Wuhan, China, April 5-8, 1999.

3. A. Blondel, Proc. of the 28th Int. Conf. on High Energy Physics, Warsaw, Poland, July 1996.

4. Z.G. Zhao, CERN COURIER Vol.38 Number 6 Sep. 1998; ’Measurement of R with BESII at BEPC’, a talk presented by Z.G. Zhao at ICHEP98, Vancouver, 1998. ”Measurement of R Between 2-5 GeV”, a talk by D. Kong, DPF99 at Los Angeles, California, Jan. 5-9, 1999; hep-ph/9903521.

5. Z.G. Zhao and F.A. Haris, CERN COURIER, Vol.39 Number 7 Sep. 1999.

6. F. A. Harris, ”Recent Charmonium Results from BES”, a talk given at Europe High Energy Phyics, Finnland July, 1998

7. F. Liu, ”Recent ψ′ Results from BES”, a talk given at Hadron 99, Beijing China, Augest 23-27, 1999.

8. J.Z. Bai et al. (BES), Phys. Lett. B (in press).

9. J.Z. Bai et al., ”A Measurement of the Mass and Full-Width of the η_{c} Meson”, to be submitted.

10. G.F. Xu, ”The analysis of the decay channels J/ψ → γγV (V = ρ, ϕ)”, a talk given at Hadron 99, Beijing China, Augest 23-27, 1999.

11. H.B. Li, ”Recent N^{∗} results from J/ψ decay”, a talk given at Hadron 99, Beijing China, Augest 23-27, 1999.

12. G. Rong, ”Direct Measurement of Br(D_{s}^{+} → \overline{K}^{∗0}X) and Br(D_{s}^{+} → K^{∗0}X), a talk given at Hadron 99, Beijing China, Augest 23-27, 1999.